OpenStudy (anonymous):

Prove e^x = sum from n=0 to infinity of (x^n)/(n!)

OpenStudy (anonymous):

I'm going to start this proof then as my question. My question is basically why we need some constant term c in bounding the Remainder of the Taylor series.

OpenStudy (anonymous):

ask*

OpenStudy (anonymous):

Let f(x) = e^x, and let \[T _{n} = \sum_{k=0}^{n}\frac{ x^k }{ k! }\]

OpenStudy (anonymous):

We can define the Remainder of the Taylor series as \[R _{n} = f(x) - T _{n}\]

OpenStudy (anonymous):

In other words, \[f(x) = T _{n} + R _{n}\]

OpenStudy (anonymous):

To show that e^x = sum from n=0 to infinity of (x^n)/(n!), it will suffice to show that \[\lim_{n \rightarrow \infty}f(x) = \lim_{n \rightarrow \infty} [T _{n} + R]\]

OpenStudy (anonymous):

And \[\lim_{n \rightarrow \infty}f(x) = f(x) = \lim_{n \rightarrow \infty}T _{n} + \lim_{n \rightarrow \infty}R _{n}\]

OpenStudy (anonymous):

Where the limit as n approaches infinity of Tn is the limit of the partial sums of the Taylor series.

OpenStudy (anonymous):

If \[\lim_{n \rightarrow \infty}R _{n} = 0\] Then f(x) = limit as n approaches infinity of Tn, which means e^x = sum from n=0 to infinity of (x^n)/(n!)

OpenStudy (anonymous):

So, we will first show that the limit as n approaches infinity of the Remainder of the Taylor series is zero.

OpenStudy (anonymous):

By Taylor's Inequality we have \[\left| R _{n} \right| \le \frac{ e^c \left| x \right|^{n+1} }{ (n + 1)! }\]

OpenStudy (anonymous):

Where c is some constant such that 0 < c < x. This is the part where I need help. Why does c have to be a constant between zero and x?

OpenStudy (anonymous):

I missed a summation symbol!

OpenStudy (anonymous):

The reason why 0 < c < x is because in the general statement of Taylor's remainder theorem, it states that there exists a value c in between x and the point you are centering your Taylor expansion around (in your case you are expanding the series around the point 0).

OpenStudy (anonymous):

By Taylor's Inequality, \[\left| R _{n} \right|=\sum_{k=n+1}^{\infty}\frac{ e^{c}\left| x \right|^{k} }{ k! }\]

OpenStudy (anonymous):

So, I think I am struggling to understand (or remember) what it means to expand the series around the point 0.

OpenStudy (anonymous):

Okay, I am looking at WolframAlpha and it is talking about Taylor Series and expanding around a point, "a".

OpenStudy (anonymous):

Does expanding around a point mean that is the starting point?

OpenStudy (anonymous):

Is it a horizontal shift?

OpenStudy (anonymous):

Ahh I see, Taylor's remainder theorem states \[|R_n| = \sum_{k=n+1}^{\infty} \frac{ e^c|x-x_0|^k }{ (n+1)! } = \frac{ e^c}{(n+1)!}(x-x_0)^{n+1}\] So in our case, x_0 is 0, mainly due to convenience.

OpenStudy (anonymous):

I understand how that works mathematically, however, conceptually I am still unsure what "expand around a point means".

OpenStudy (anonymous):

@ganeshie8 Do you know what "expand around a point" means conceptually in relation to a Taylor Series Expansion?

OpenStudy (anonymous):

Okay, so the point chosen will affect the radius of convergence for the series.

OpenStudy (anonymous):

When you expand the taylor series around an arbitrary point, you are basically forming a better and better approximation to your original function (at that specific location) by higher degree polynomials. And as that higher degree polynomial goes towards an infinite degree (taking n to infinity), your function (at that specific point) actually becomes the taylors series expansion

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I think I can finish the proof from here.

OpenStudy (anonymous):

Cool : )

OpenStudy (anonymous):

So, \[\left| R _{n} \right|\le \sum_{k=n+1}^{\infty}\frac{ e ^{c}\left| x \right|^{k} }{ k! } = e ^{c}\sum_{k=n+1}^{\infty}\frac{ \left| x \right|^{k} }{ k! } = e ^{c}*(0) = 0\]

OpenStudy (anonymous):

Therefore, \[\left| R _{n} \right|\le0 \] Which implies that the Remainder is equal to zero by the squeeze theorem. Thus f(x) = limit as n approaches infinity of the Partial Sum of the Taylor Series, and so e^x = the sum from n=0 to infinity of (x^n)/n!

OpenStudy (anonymous):

It seems obvious but I'd still be wary of the use of (without explicit proof) \[\lim_{k \rightarrow \infty} \sum_{k=n+1}^{\infty} \frac{|x|^k}{k!} = 0\] All else looks good. Nice job : )

OpenStudy (anonymous):

Thanks for your help domu! Great point, I'll go ahead and prove that to myself.

OpenStudy (anonymous):

By the Ratio Test, \[\frac{ \frac{ \left| x \right|^{k+1} }{ (k+1)! } }{ \frac{ \left| x \right|^{k} }{ (k)! } }=\frac{ \left| x \right|^{k+1} k! }{ \left| x \right|^k (k+1)! }=\frac{ \left| x \right| }{ k+1 } = 0\]

OpenStudy (anonymous):

This is true for all real values of x since k is approaching infinity. Thus, by the Ratio Test, this series converges to zero and has an infinite radius of convergence.

OpenStudy (anonymous):

So at this point I think I'm done. Feel free to ask me any questions or give me any comments.

OpenStudy (experimentx):

define \[ e = \lim_{n \to \infty } \left( 1 + \frac 1 n\right)^n \] and expand it nominally show that \[ \lim_{n\to \infty } \left | \left( 1 + \frac 1 n\right)^n - \sum_{k=1}^\infty \frac 1 {k!} \right |<\epsilon \]